# YOLO26: Key Architectural Enhancements and Performance Benchmarking for Real-Time Object Detection

### v2

@@ 1 Introduction
Object detection has emerged as one of the most critical tasks in computer vision, enabling machines to localize and classify multiple objects within an image or video stream zhao2019object ; zou2023object . From autonomous driving and robotics to surveillance, medical imaging, agriculture, and smart manufacturing, real-time object detection algorithms serve as the backbone of artificial intelligence (AI) applications rana2024artificial ; khan2025objectdetection . Among these algorithms, the You Only Look Once (YOLO) family has established itself as the most influential series of models for real-time object detection, combining accuracy with unprecedented inference speed sapkota2025yolo ; sapkota2025rf ; sapkota2024comparing ; sapkota2024comparing . Since its introduction in 2016, YOLO has evolved through numerous architectural revisions, each addressing limitations of its predecessors while integrating cutting-edge advances in neural network design, loss functions, and deployment efficiency sapkota2025yolo . The release of YOLO26 in September 2025 represents the latest milestone in this evolutionary trajectory, introducing architectural simplifications, novel optimizers, and enhanced edge deployment capabilities designed for low-power devices.
@@
YOLOv1 (2015) redmon2016you First unified single-stage object detector (one network for bounding boxes + class probabilities). Object Detection, Classification Darknet
YOLOv2 (2016) redmon2017yolo9000 Multi-scale training introduced; anchor box dimension clustering for improved prior boxes (YOLO9000 joint detection/classification). Object Detection, Classification Darknet
YOLOv3 (2018) redmon2018yolov3 Deeper Darknet-53 backbone with residual connections; added SPP module and multi-scale feature fusion for small object detection. Object Detection, Multi-scale Detection Darknet
YOLOv4 (2020) bochkovskiy2020yolov4 Mish activation function adopted; CSPDarknet-53 backbone (Cross-Stage Partial networks) for enhanced feature reuse. Object Detection, Object Tracking Darknet
@@
YOLOv6 (2022) li2022yolov6 EfficientRep backbone with embedded self-attention; introduced anchor-free object detection mode for efficiency. Object Detection, Instance Segmentation PyTorch
YOLOv7 (2022) wang2023yolov7 Extended ELAN (E-ELAN) backbone with model re-parameterization; integrated transformer-based modules for broader tasks (e.g. tracking). Object Detection, Object Tracking, Instance Segmentation PyTorch
@@
YOLOv9 (2024) wang2024yolov9 Introduced Programmable Gradient Information (PGI) for selective learning; proposed G-ELAN (an enhanced ELAN architecture) for improved feature extraction. Object Detection, Instance Segmentation PyTorch
YOLOv10 (2024) wang2024yolov10 Achieved end-to-end NMS-free detection via a consistent dual-assignment training strategy (removing post-processing). Object Detection PyTorch
@@
YOLOv12 (2025) tian2025yolov12 Attention-centric architecture: introduced an efficient area attention module (global self-attention with low complexity) and Residual ELAN (R-ELAN) blocks to improve feature aggregation, achieving transformer-level accuracy at YOLO speeds. Object Detection PyTorch
YOLOv13 (2025) lei2025yolov13 Hypergraph-based Adaptive Correlation Enhancement (HyperACE) module to capture global high-order feature interactions; Full-Pipeline Aggregation-Distribution (FullPAD) scheme for enhanced feature flow throughout the network; utilized depthwise-separable convolutions to reduce complexity. Object Detection PyTorch
@@
The YOLO framework was first proposed by Joseph Redmon and colleagues in 2016, introducing a paradigm shift in object detection redmon2016you . Unlike traditional two-stage detectors such as R-CNN he2017mask and Faster R-CNN ren2016faster , which separated region proposal from classification, YOLO formulated detection as a single regression problem diwan2023object . By directly predicting bounding boxes and class probabilities in one forward pass through a convolutional neural network (CNN), YOLO achieved real-time speeds while maintaining competitive accuracy ali2024yolo ; diwan2023object . This efficiency made YOLOv1 highly attractive for applications where latency was a critical factor, including robotics, autonomous navigation, and live video analytics. Subsequent versions YOLOv2 (2017) redmon2017yolo9000 and YOLOv3 (2018) redmon2018yolov3 significantly improved accuracy while retaining real-time performance. YOLOv2 introduced batch normalization, anchor boxes, and multi-scale training, which increased robustness across varying object sizes. YOLOv3 leveraged a deeper architecture based on Darknet-53, along with multi-scale feature maps for better small-object detection. These enhancements made YOLOv3 the de facto standard for academic and industrial applications for several years apostolidis2025delving ; sapkota2025yolo ; edozie2025comprehensive .
@@
As the demand for higher accuracy grew, especially in challenging domains such as aerial imagery, agriculture, and medical analysis, YOLO models diversified into more advanced architectures. YOLOv4 (2020) bochkovskiy2020yolov4 introduced Cross-Stage Partial Networks (CSPNet), improved activation functions like Mish, and advanced training strategies including mosaic data augmentation and CIoU loss. YOLOv5 (Ultralytics, 2020), though unofficial, gained immense popularity due to its PyTorch implementation, extensive community support, and simplified deployment across diverse platforms. YOLOv5 also brought modularity, making it easier to adapt for segmentation, classification, and edge applications. Further developments included YOLOv6li2022yolov6 and YOLOv7 wang2023yolov7 (2022), which integrated advanced optimization techniques, parameter-efficient modules, and transformer-inspired blocks. These iterations pushed YOLO closer to state-of-the-art (SoTA) accuracy benchmarks while retaining a focus on real-time inference. The YOLO ecosystem, by this point, had firmly established itself as the leading family of models in object detection research and deployment.
@@
Ultralytics, the primary maintainer of modern YOLO releases, redefined the framework with YOLOv8 (2023) sohan2024review . YOLOv8 featured a decoupled detection head, anchor-free predictions, and refined training strategies, resulting in substantial improvements in both accuracy and deployment versatility farooq2024improved . It was widely adopted in industry due to its clean Python API, compatibility with TensorRT, CoreML, and ONNX, and availability of variants optimized for speed versus accuracy trade-offs (nano, small, medium, large, and extra-large). YOLOv9 wang2024yolov9 , YOLOv10 wang2024yolov10 , and YOLO11 followed in rapid succession, each iteration pushing the boundaries of architecture and performance. YOLOv9 introduced GELAN (Generalized Efficient Layer Aggregation Network) and Progressive Distillation, combining efficiency with higher representational capacity. YOLOv10 focused on balancing accuracy and inference latency with hybrid task-aligned assignments. YOLOv11 further refined Ultralytics’ vision, offering higher efficiency on GPUs while maintaining strong small-object performance sapkota2025yolo . Together, these models cemented Ultralytics’ reputation for producing production-ready YOLO releases tailored to modern deployment pipelines.
@@
Following YOLO11, alternative versions YOLOv12tian2025yolov12 and YOLOv13 lei2025yolov13 introduced attention-centric designs and advanced architectural components that sought to maximize accuracy across diverse datasets. These models explored multi-head self-attention, improved multi-scale fusion, and stronger training regularization strategies. While they offered strong benchmarks, they retained reliance on Non-Maximum Suppression (NMS) and Distribution Focal Loss (DFL), which introduced latency overhead and export challenges, especially for low-power devices. The limitations of NMS-based post-processing and complex loss formulations motivated the development of YOLO26 (Ultralytics YOLO26 Official Source). By September 2025, at the YOLO Vision 2025 event in London, Ultralytics unveiled YOLO26 as a next-generation model optimized for edge computing, robotics, and mobile AI.
@@ 3 Benchmarking and Comparative Analysis
In the case of YOLO26, a series of rigorous benchmarks were conducted to assess its performance in comparison to both its YOLO predecessors and alternative state-of-the-art architectures. Figure 4 presents a consolidated view of this evaluation, plotting COCO mAP(50–95) against latency (ms per image) on an NVIDIA T4 GPU with TensorRT FP16 optimization. The inclusion of competing architectures such as YOLOv10, RT-DETR, RT-DETRv2, RT-DETRv3, and DEIM provides a comprehensive landscape of recent advancements in real-time detection. From the figure, YOLO26 demonstrates a distinctive positioning: it maintains high accuracy levels that rival transformer-based models like RT-DETRv3, while significantly outperforming them in terms of inference speed. For instance, YOLO26-m and YOLO26-l achieve competitive mAP scores above 51% and 53%, respectively, but at a substantially reduced latency, underscoring the benefits of its NMS-free architecture and lightweight regression head.
Figure 4: Performance benchmarking of YOLO26 compared with YOLOv10, RT-DETR, RT-DETRv2, RT-DETRv3, and DEIM on the COCO dataset. The plot shows COCO mAP(50–95) versus latency (ms per image) measured on an NVIDIA T4 GPU using TensorRT FP16 inference. YOLO26 demonstrates superior balance between accuracy and efficiency, achieving competitive detection performance while significantly reducing latency, thereby highlighting its suitability for real-time edge and resource-constrained deployments.
This balance between accuracy and speed is particularly relevant for edge deployments, where maintaining real-time throughput is as important as ensuring reliable detection quality. Compared with YOLOv10, YOLO26 consistently achieves lower latency across model scales, with speedups of up to 43% observed for CPU-bound inference, while preserving or improving accuracy through its ProgLoss and STAL mechanisms. When compared to DEIM and the RT-DETR series, which rely heavily on transformer encoders and decoders, YOLO26’s simplified backbone and MuSGD-driven training pipeline enable faster convergence and leaner inference without compromising small-object recognition. The plot in Figure 4 clearly illustrates these distinctions: while RT-DETRv3 excels in large-scale accuracy benchmarks, its latency profile remains less favorable than YOLO26, reinforcing YOLO26’s edge-centric design philosophy. Furthermore, the benchmarking analysis highlights YOLO26’s robustness in balancing the accuracy–latency curve, situating it as a versatile detector suitable for both high-throughput server applications and resource-constrained devices. This comparative evidence substantiates the claim that YOLO26 is not merely an incremental update but a paradigm shift in the YOLO lineage, successfully bridging the gap between the efficiency-first philosophy of earlier YOLO models and the accuracy-driven orientation of transformer-based detectors. Ultimately, the benchmarking results demonstrate that YOLO26 offers a compelling deployment advantage, particularly in real-world environments requiring reliable performance under stringent latency constraints.

### v3

@@ 1 Introduction
Object detection has emerged as one of the most critical tasks in computer vision, enabling machines to localize and classify multiple objects within an image or video stream Zhao et al. (2019); Zou et al. (2023). From autonomous driving and robotics to surveillance, medical imaging, agriculture, and smart manufacturing, real-time object detection algorithms serve as the backbone of artificial intelligence (AI) applications Rana and others (2024); Khan et al. (2025). Among these algorithms, the You Only Look Once (YOLO) family has established itself as the most influential series of models for real-time object detection, combining accuracy with unprecedented inference speed Sapkota et al. (2025b, a, 2024, 2024). Since its introduction in 2016, YOLO has evolved through numerous architectural revisions, each addressing limitations of its predecessors while integrating cutting-edge advances in neural network design, loss functions, and deployment efficiency Sapkota et al. (2025b).
The release of YOLO26 in September 2025 marks the newest milestone in the YOLO lineage, shifting the design emphasis from incremental architectural complexity toward deployment-oriented simplification—most notably through streamlined regression, end-to-end prediction behavior, and training-time refinements enabled by novel optimization. This edge-first philosophy is reflected in the comparative accuracy–latency trends shown in Fig. 1a, where Ultralytics reports YOLO26’s COCO mAP(50-95) versus latency performance (T4, TensorRT10, FP16) against a broad set of prior YOLO variants (YOLO11, YOLOv10, YOLOv9, YOLOv8, YOLOv7, YOLOv6-3.0, YOLOv5) as well as competitive real-time detectors (PP-YOLOE+, DAMO-YOLO, and RTMDet). Complementing this, Fig. 1b positions YOLO26 on the same COCO mAP(50-95) versus end-to-end latency axis against transformer-style real-time baselines (YOLOv10 and the RT-DETR family), underscoring that YOLO26 aims to retain high detection quality while reducing overall pipeline delay, an especially relevant trade-off for low-power and latency-sensitive edge devices.
Figure 1: Performance comparison of YOLO26 under TensorRT FP16 on an NVIDIA T4 GPU (Source Link). (a) COCO mAP(50–95) versus inference latency (ms/image), comparing YOLO26 with earlier YOLO versions and other real-time detectors, highlighting its improved accuracy–speed trade-off. (b) COCO mAP(50–95) versus end-to-end latency, comparing YOLO26 with YOLOv10 and RT-DETR variants, illustrating its advantage in overall pipeline efficiency.
@@
YOLOv1 (2015) Redmon et al. (2016) First unified single-stage object detector (one network for bounding boxes + class probabilities). Object Detection, Classification Darknet
YOLOv2 (2016) Redmon and Farhadi (2017) Multi-scale training introduced; anchor box dimension clustering for improved prior boxes (YOLO9000 joint detection/classification). Object Detection, Classification Darknet
YOLOv3 (2018) Redmon and Farhadi (2018) Deeper Darknet-53 backbone with residual connections; added SPP module and multi-scale feature fusion for small object detection. Object Detection, Multi-scale Detection Darknet
YOLOv4 (2020) Bochkovskiy et al. (2020) Mish activation function adopted; CSPDarknet-53 backbone (Cross-Stage Partial networks) for enhanced feature reuse. Object Detection, Object Tracking Darknet
@@
YOLOv6 (2022) Li et al. (2022a) EfficientRep backbone with embedded self-attention; introduced anchor-free object detection mode for efficiency. Object Detection, Instance Segmentation PyTorch
YOLOv7 (2022) Wang et al. (2023) Extended ELAN (E-ELAN) backbone with model re-parameterization; integrated transformer-based modules for broader tasks (e.g. tracking). Object Detection, Object Tracking, Instance Segmentation PyTorch
@@
YOLOv9 (2024) Wang et al. (2024b) Introduced Programmable Gradient Information (PGI) for selective learning; proposed G-ELAN (an enhanced ELAN architecture) for improved feature extraction. Object Detection, Instance Segmentation PyTorch
YOLOv10 (2024) Wang et al. (2024a) Achieved end-to-end NMS-free detection via a consistent dual-assignment training strategy (removing post-processing). Object Detection PyTorch
@@
YOLOv12 (2025) Tian et al. (2025b) Attention-centric architecture: introduced an efficient area attention module (global self-attention with low complexity) and Residual ELAN (R-ELAN) blocks to improve feature aggregation, achieving transformer-level accuracy at YOLO speeds. Object Detection PyTorch
YOLOv13 (2025) Lei et al. (2025) Hypergraph-based Adaptive Correlation Enhancement (HyperACE) module to capture global high-order feature interactions; Full-Pipeline Aggregation-Distribution (FullPAD) scheme for enhanced feature flow throughout the network; utilized depthwise-separable convolutions to reduce complexity. Object Detection PyTorch
@@
The YOLO framework was first proposed by Joseph Redmon and colleagues in 2016, introducing a paradigm shift in object detection Redmon et al. (2016). Unlike traditional two-stage detectors such as R-CNN He et al. (2017) and Faster R-CNN Ren et al. (2016), which separated region proposal from classification, YOLO formulated detection as a single regression problem Diwan et al. (2023). By directly predicting bounding boxes and class probabilities in one forward pass through a convolutional neural network (CNN), YOLO achieved real-time speeds while maintaining competitive accuracy Ali and Zhang (2024); Diwan et al. (2023). This efficiency made YOLOv1 highly attractive for applications where latency was a critical factor, including robotics, autonomous navigation, and live video analytics. Subsequent versions YOLOv2 (2017) Redmon and Farhadi (2017)and YOLOv3 (2018) Redmon and Farhadi (2018) significantly improved accuracy while retaining real-time performance. YOLOv2 introduced batch normalization, anchor boxes, and multi-scale training, which increased robustness across varying object sizes. YOLOv3 leveraged a deeper architecture based on Darknet-53, along with multi-scale feature maps for better small-object detection. These enhancements made YOLOv3 the de facto standard for academic and industrial applications for several years Apostolidis and Papakostas (2025); Sapkota et al. (2025b); Edozie et al. (2025).
@@
As the demand for higher accuracy grew, especially in challenging domains such as aerial imagery, agriculture, and medical analysis, YOLO models diversified into more advanced architectures. YOLOv4 (2020) Bochkovskiy et al. (2020) introduced Cross-Stage Partial Networks (CSPNet), improved activation functions like Mish, and advanced training strategies including mosaic data augmentation and CIoU loss. YOLOv5 (Ultralytics, 2020), though unofficial, gained immense popularity due to its PyTorch implementation, extensive community support, and simplified deployment across diverse platforms. YOLOv5 also brought modularity, making it easier to adapt for segmentation, classification, and edge applications. Further developments included YOLOv6Li et al. (2022a) and YOLOv7 Wang et al. (2023) (2022), which integrated advanced optimization techniques, parameter-efficient modules, and transformer-inspired blocks. These iterations pushed YOLO closer to state-of-the-art (SoTA) accuracy benchmarks while retaining a focus on real-time inference. The YOLO ecosystem, by this point, had firmly established itself as the leading family of models in object detection research and deployment.
@@
Ultralytics, the primary maintainer of modern YOLO releases, redefined the framework with YOLOv8 (2023) Sohan et al. (2024). YOLOv8 featured a decoupled detection head, anchor-free predictions, and refined training strategies, resulting in substantial improvements in both accuracy and deployment versatility Farooq et al. (2024). It was widely adopted in industry due to its clean Python API, compatibility with TensorRT, CoreML, and ONNX, and availability of variants optimized for speed versus accuracy trade-offs (nano, small, medium, large, and extra-large). YOLOv9 Wang et al. (2024b), YOLOv10 Wang et al. (2024a), and YOLO11 followed in rapid succession, each iteration pushing the boundaries of architecture and performance. YOLOv9 introduced GELAN (Generalized Efficient Layer Aggregation Network) and Progressive Distillation, combining efficiency with higher representational capacity. YOLOv10 focused on balancing accuracy and inference latency with hybrid task-aligned assignments. YOLOv11 further refined Ultralytics’ vision, offering higher efficiency on GPUs while maintaining strong small-object performance Sapkota et al. (2025b). Together, these models cemented Ultralytics’ reputation for producing production-ready YOLO releases tailored to modern deployment pipelines.
@@
Following YOLO11, alternative versions YOLOv12Tian et al. (2025b) and YOLOv13 Lei et al. (2025) introduced attention-centric designs and advanced architectural components that sought to maximize accuracy across diverse datasets. These models explored multi-head self-attention, improved multi-scale fusion, and stronger training regularization strategies. While they offered strong benchmarks, they retained reliance on Non-Maximum Suppression (NMS) and Distribution Focal Loss (DFL), which introduced latency overhead and export challenges, especially for low-power devices. The limitations of NMS-based post-processing and complex loss formulations motivated the development of YOLO26 (Ultralytics YOLO26 Official Source). By September 2025, at the YOLO Vision 2025 event in London, Ultralytics unveiled YOLO26 as a next-generation model optimized for edge computing, robotics, and mobile AI.
@@ 3 Benchmarking and Comparative Analysis
3.1 Detection and Segmentation Performance Metrics
As summarized in the Detection results in Table 2, YOLO26 consistently improves COCO mAP(50–95) as model scale increases, while maintaining predictable and low inference latency across CPU (ONNX) and GPU (TensorRT) runtimes. In particular, YOLO26-m and YOLO26-l achieve strong detection accuracy above 53% and 55% mAP, respectively, at substantially lower latency than transformer-based alternatives, reflecting the benefits of its NMS-free inference path and simplified regression design. The same table further highlights favorable scaling behavior in parameters and FLOPs, reinforcing YOLO26’s efficiency across deployment targets.
Beyond detection, the Segmentation results in Table 2 demonstrate that YOLO26 retains these advantages in multi-task settings. Across nano to extra-large variants, YOLO26-seg models deliver competitive box and mask mAP while preserving manageable computational cost and real-time throughput, even under end-to-end evaluation. When contrasted with architectures such as YOLOv10 and RT-DETR variants, which rely on heavier transformer encoders, YOLO26 exhibits a more balanced accuracy–latency profile, particularly for edge and CPU-bound inference. Taken together, the detection and segmentation benchmarks in Table 2 show that YOLO26 is not merely an incremental refinement, but a deployment-oriented evolution of the YOLO family, effectively bridging efficiency-focused design and high-accuracy real-time perception under stringent latency constraints.
Table 2: Ultralytics YOLO26 performance metrics (640 px). Detection (top) and instance segmentation (bottom) results report COCO validation accuracy, end-to-end (e2e) scores where applicable, and speed on CPU (ONNX) and NVIDIA T4 (TensorRT10 FP16), along with model size (params) and compute (FLOPs).
Detection
Model Size mAPval mAPval Speed Speed Params FLOPs
(px) 50–95 50–95 (e2e) CPU ONNX (ms) T4 TRT10 (ms) (M) (B)
YOLO26n 640 40.9 40.1 $38.9\pm 0.7$ $1.7\pm 0.0$ 2.4 5.4
YOLO26s 640 48.6 47.8 $87.2\pm 0.9$ $2.5\pm 0.0$ 9.5 20.7
YOLO26m 640 53.1 52.5 $220.0\pm 1.4$ $4.7\pm 0.1$ 20.4 68.2
YOLO26l 640 55.0 54.4 $286.2\pm 2.0$ $6.2\pm 0.2$ 24.8 86.4
YOLO26x 640 57.5 56.9 $525.8\pm 4.0$ $11.8\pm 0.2$ 55.7 193.9
Instance Segmentation
Model Size mAPbox mAPmask Speed Speed Params FLOPs
(px) 50–95 (e2e) 50–95 (e2e) CPU ONNX (ms) T4 TRT10 (ms) (M) (B)
YOLO26n-seg 640 39.6 33.9 $53.3\pm 0.5$ $2.1\pm 0.0$ 2.7 9.1
YOLO26s-seg 640 47.3 40.0 $118.4\pm 0.9$ $3.3\pm 0.0$ 10.4 34.2
YOLO26m-seg 640 52.5 44.1 $328.2\pm 2.4$ $6.7\pm 0.1$ 23.6 121.5
YOLO26l-seg 640 54.4 45.5 $387.0\pm 3.7$ $8.0\pm 0.1$ 28.0 139.8
YOLO26x-seg 640 56.5 47.0 $787.0\pm 6.8$ $16.4\pm 0.1$ 62.8 313.5
3.2 Classification performance metrics (ImageNet)
Table 3 summarizes the ImageNet classification performance of YOLO26 across model scales. As model capacity increases from YOLO26n to YOLO26x, Top-1 accuracy improves steadily from 71.4% to 79.9%, while maintaining strong Top-5 accuracy above 90% for all variants. Importantly, this accuracy scaling is achieved with predictable latency growth, as TensorRT FP16 inference remains below 4 ms even for the largest model. The compact FLOPs and parameter counts reported in Table 3 highlight that YOLO26 classification heads preserve efficiency, making them well suited for real-time image recognition on edge and embedded platforms.
Table 3: YOLO26 image classification performance on ImageNet at 224 px resolution. Results report Top-1/Top-5 accuracy, inference speed on CPU (ONNX) and NVIDIA T4 (TensorRT10 FP16), along with model size and FLOPs.
Model Size Acc Acc Speed Speed Params FLOPs
(px) Top-1 Top-5 CPU ONNX (ms) T4 TRT10 (ms) (M) (B @224)
YOLO26n-cls 224 71.4 90.1 $5.0\pm 0.3$ $1.1\pm 0.0$ 2.8 0.5
YOLO26s-cls 224 76.0 92.9 $7.9\pm 0.2$ $1.3\pm 0.0$ 6.7 1.6
YOLO26m-cls 224 78.1 94.2 $17.2\pm 0.4$ $2.0\pm 0.0$ 11.6 4.9
YOLO26l-cls 224 79.0 94.6 $23.2\pm 0.3$ $2.8\pm 0.0$ 14.1 6.2
YOLO26x-cls 224 79.9 95.0 $41.4\pm 0.9$ $3.8\pm 0.0$ 29.6 13.6
3.3 Pose Performance Metrics (COCO)
Table 4 presents the pose estimation performance of YOLO26 on the COCO dataset. Across model scales, YOLO26 exhibits consistent gains in mAPpose, increasing from 57.2% for the nano variant to 71.6% for the extra-large model under end-to-end evaluation. This accuracy improvement is accompanied by predictable scaling in latency and computational cost, while maintaining real-time inference on GPU and near-real-time performance on CPU. The results in Table 4 demonstrate that YOLO26 effectively extends its efficiency-oriented design to pose estimation, making it suitable for real-time human and object keypoint analysis on both edge and server platforms.
Table 4: YOLO26 pose estimation performance on the COCO dataset at 640 px resolution. Results report end-to-end (e2e) pose accuracy, inference speed on CPU (ONNX) and NVIDIA T4 (TensorRT10 FP16), along with model size and FLOPs.
Model Size mAPpose mAPpose Speed Speed Params FLOPs
(px) 50–95 (e2e) 50 (e2e) CPU ONNX (ms) T4 TRT10 (ms) (M) (B)
YOLO26n-pose 640 57.2 83.3 $40.3\pm 0.5$ $1.8\pm 0.0$ 2.9 7.5
YOLO26s-pose 640 63.0 86.6 $85.3\pm 0.9$ $2.7\pm 0.0$ 10.4 23.9
YOLO26m-pose 640 68.8 89.6 $218.0\pm 1.5$ $5.0\pm 0.1$ 21.5 73.1
YOLO26l-pose 640 70.4 90.5 $275.4\pm 2.4$ $6.5\pm 0.1$ 25.9 91.3
YOLO26x-pose 640 71.6 91.6 $565.4\pm 3.0$ $12.2\pm 0.2$ 57.6 201.7
3.4 Oriented Object Detection (OBB) Performance on DOTA v1
[…]
Table 5: YOLO26 oriented object detection (OBB) performance on the DOTA v1 dataset at 1024 px resolution. Results report end-to-end (e2e) test accuracy, inference speed on CPU (ONNX) and NVIDIA T4 (TensorRT10 FP16), along with model size and FLOPs.
Model Size mAPtest mAPtest Speed Speed Params FLOPs
(px) 50–95 (e2e) 50 (e2e) CPU ONNX (ms) T4 TRT10 (ms) (M) (B)
YOLO26n-obb 1024 52.4 78.9 $97.7\pm 0.9$ $2.8\pm 0.0$ 2.5 14.0
YOLO26s-obb 1024 54.8 80.9 $218.0\pm 1.4$ $4.9\pm 0.1$ 9.8 55.1
YOLO26m-obb 1024 55.3 81.0 $579.2\pm 3.8$ $10.2\pm 0.3$ 21.2 183.3
YOLO26l-obb 1024 56.2 81.6 $735.6\pm 3.1$ $13.0\pm 0.2$ 25.6 230.0
YOLO26x-obb 1024 56.7 81.7 $1485.7\pm 11.5$ $30.5\pm 0.9$ 57.6 516.5